OpenStudy (kitkat16):

Anyone know if this is correct and why? In a tire factory, a machine is responsible for molding the tire. It has a failure rate of 0.2%. If 1000 tires are produced in a day, and 6 are faulty. What is the difference between the experimental probability and the theoretical probability? 0.4%

OpenStudy (phi):

theoretically 0.2% times 1000 or 0.2/100 * 1000 = 0.2*10 = 2 observed: 6 difference is 6-2 = 4 yes, 4 is correct.

OpenStudy (kitkat16):

Thanks!!!